# ORCID: 0009-0007-8992-9682

### v2

@@ 5 Experiment Results
A Larger model size does not always guarantee better robustness or pass rate across all programming languages. As shown in Table 8 and Table 9, in general, we can observe a robustness or pass rate improvement with the increase of model sizes. For instance, for CodeGen in the worst case scenario, the $RD_{5}@1$ of JavaScript increases from 0.07 to 0.08 for 2B and 6B, respectively. On the contrary, for Java, we also observe that there is a drop of 0.02 of the $RD_{5}@1$. The results show that increasing the model size does not always guarantee better robustness and highlight the need for more comprehensive experiments on different programming languages.
CodeGen has a relatively higher pass@k than Incoder across different languages. By checking the Nominal metric in Table 8 and Table 9, with the same parameter size (e.g., 6B), CodeGen usually has a better performance than that of Incoder. For example, for Java, CodeGen has a pass@k of 0.16 compared to 0.11 produced by Incoder. This might be due to the different sizes of tokens used for the training, as CodeGen has been trained on more than 1000 billion tokens and Incoder has only been trained on about 60 billion tokens. The other reason may come from the distribution of the training data. Although the training data of Incoder contains 28 languages, which is much more diverse than that of CodeGen (i.e., trained in six programming languages), the majority of the data is Python and JavaScript. This may also explain why Incoder only has the worst pass@k of 0.02 on the Java coding problems.
@@
Table 8: Performance summary of perturbations against code generation models using HumanEval-X test cases and worst case passes (5 out of 5). Up arrows indicate the extent of performance improvement, while down arrows indicate the extent of performance drop.
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
@@
RP5@1$\uparrow$ 0.0 0.04 0.04 0.04 0.07 0.07 0.08 0.07 0.07 0.05 0.08 0.07
RD5@1$\downarrow$ 0.88 0.57 0.49 0.64 0.46 0.47 0.4 0.44 0.48 0.69 0.4 0.56
RR5@1$\downarrow$ 0.02 0.07 0.05 0.08 0.06 0.07 0.05 0.06 0.07 0.11 0.06 0.11
Syntax Nominal (Partial)$\uparrow$ 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44
RP5@1$\uparrow$ 0.23 0.03 0.23 0.27 0.05 0.27 0.33 0.07 0.2 0.29 0.09 0.23
RD5@1$\downarrow$ 0.3 0.73 0.39 0.38 0.68 0.39 0.41 0.45 0.49 0.46 0.58 0.47
RR5@1$\downarrow$ 0.21 0.08 0.16 0.24 0.14 0.17 0.26 0.08 0.2 0.25 0.14 0.24
Format Nominal (Partial)$\uparrow$ 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44
RP5@1$\uparrow$ 0.2 0.04 0.26 0.28 0.07 0.3 0.45 0.1 0.27 0.42 0.12 0.27
RD5@1$\downarrow$ 0.38 0.64 0.33 0.36 0.57 0.32 0.2 0.22 0.31 0.21 0.44 0.39
RR5@1$\downarrow$ 0.16 0.07 0.14 0.19 0.11 0.15 0.13 0.04 0.16 0.13 0.1 0.21
Table 9: Performance summary of perturbations using partial cases using HuamEval-X test cases and partial passes (3 out of 5)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
@@
RP5@1$\uparrow$ 0.03 0.08 0.08 0.07 0.1 0.12 0.12 0.12 0.11 0.11 0.12 0.15
RD5@1$\downarrow$ -0.04 0.19 -0.01 0.37 0.15 0.1 0.06 0.05 0.11 0.28 0.12 0.12
RR5@1$\downarrow$ 0.02 0.05 0.04 0.07 0.05 0.05 0.05 0.05 0.05 0.09 0.04 0.07
Syntax Nominal (Partial)$\uparrow$ 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44
RP5@1$\uparrow$ 0.27 0.07 0.28 0.36 0.11 0.32 0.4 0.11 0.28 0.39 0.14 0.33
RD5@1$\downarrow$ 0.18 0.37 0.26 0.19 0.36 0.27 0.29 0.13 0.28 0.27 0.35 0.24
RR5@1$\downarrow$ 0.2 0.08 0.11 0.2 0.17 0.13 0.21 0.09 0.16 0.18 0.13 0.17
Format Nominal (Partial)$\uparrow$ 0.33 0.1 0.38 0.45 0.17 0.44 0.56 0.13 0.39 0.54 0.22 0.44
RP5@1$\uparrow$ 0.32 0.07 0.32 0.39 0.13 0.41 0.52 0.14 0.36 0.53 0.19 0.35
RD5@1$\downarrow$ 0.03 0.29 0.18 0.12 0.21 0.06 0.07 -0.07 0.09 0.01 0.12 0.2
RR5@1$\downarrow$ 0.11 0.05 0.09 0.14 0.1 0.07 0.08 0.04 0.12 0.05 0.06 0.15
RQ2: What are the major factors that contribute to the difference in the robustness of LLMs on different programming languages?
Approach: To answer this research question, we first find out the error rates for the nominal, partial, and each category of perturbations. We categorize the run statuses as passed, assertion error, compilation error, timeout, and runtime error. To calculate these ratios, we quantify the total occurrence of each category and divide it by the overall number of instances, thereby providing insights into the distribution of these outcomes. Subsequently, we engage in a manual analysis of select samples from each programming language, delving into the underlying causes of failures and variations in robustness across different languages.
Results: The resultant error rates are presented in Table 10. From the table, it is observed that nominal prompts exhibit similar pass rates across the three languages. Furthermore, the pass rates for perturbations stemming from nominal prompts display analogous patterns. On the other hand, notable differences emerge in the error rates associated with partial data. In partial cases, C++ shows a lower pass rate compared to Java and JavaScript. Notably, C++ yields a significantly higher count of assertion errors in comparison to the other two languages, suggesting that C++ fares less favorably in code completion tasks when contrasted with Java and JavaScript.
Table 10: Error Percentages for code generation models
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Type Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal Passed 2.44 10.37 7.93 10.98 12.2 12.8 12.8 12.2 12.8 15.85 14.02 16.46
Assertion 31.1 60.37 55.49 55.49 62.8 63.41 69.51 68.29 60.98 66.46 67.07 48.78
Compilation 58.54 29.27 35.98 26.83 22.56 23.17 9.15 18.29 25.61 5.49 17.07 33.54
Timeout 1.83 0.0 0.61 0.61 0.61 0.61 0.61 0.61 0.61 1.22 1.22 1.22
Runtime 6.1 0.0 0.0 6.1 1.83 0.0 7.93 0.61 0.0 10.98 0.61 0.0
DocString Passed 2.9 8.57 6.8 8.17 11.37 11.27 12.43 12.87 10.51 13.0 13.9 12.88
Assertion 33.3 55.95 49.07 51.13 60.74 62.72 65.2 63.46 49.71 57.46 65.6 42.85
Compilation 57.32 34.54 42.21 32.07 26.67 25.49 14.16 22.45 39.04 21.95 18.7 43.57
Timeout 0.95 0.67 1.91 1.04 0.44 0.52 1.0 0.78 0.74 1.37 1.11 0.7
Runtime 5.52 0.27 0.0 7.59 0.78 0.0 7.22 0.44 0.0 6.22 0.7 0.0
Function Passed 2.85 9.13 7.89 8.6 11.43 12.2 12.56 12.63 12.01 12.46 13.34 14.74
Assertion 34.47 59.09 54.55 55.06 64.39 64.43 68.43 67.3 56.1 61.4 69.58 47.36
Compilation 55.63 30.71 35.71 27.05 22.56 22.97 10.93 18.87 31.04 18.9 14.77 37.05
Timeout 1.06 0.89 1.85 0.83 0.52 0.41 1.02 0.71 0.85 0.77 1.36 0.85
Runtime 6.0 0.17 0.0 8.46 1.1 0.0 7.05 0.49 0.0 6.46 0.94 0.0
Partial Passed 32.93 10.37 38.41 44.51 17.07 43.9 56.1 12.8 39.02 53.66 21.95 43.9
Assertion 31.71 62.2 48.78 42.07 63.41 44.51 35.98 67.07 43.9 38.41 65.24 37.8
Compilation 30.49 26.22 12.8 9.15 17.68 11.59 5.49 17.07 17.07 4.27 10.98 17.68
Timeout 1.83 0.61 0.0 1.83 0.61 0.0 0.61 1.83 0.0 0.0 0.61 0.61
Runtime 3.05 0.61 0.0 2.44 1.22 0.0 1.83 1.22 0.0 3.66 1.22 0.0
Syntax Passed 27.66 7.95 28.64 36.67 12.32 32.78 40.59 12.09 29.07 39.09 14.8 33.46
Assertion 25.79 45.79 43.15 28.39 50.63 41.32 32.93 56.38 40.08 28.8 55.53 39.04
Compilation 41.97 44.92 27.28 31.4 35.69 24.84 22.28 30.18 29.94 28.05 28.17 26.99
Timeout 1.73 1.1 0.93 0.59 1.14 1.06 0.57 1.08 0.91 0.71 1.28 0.51
Runtime 2.85 0.24 0.0 2.95 0.22 0.0 3.64 0.26 0.0 3.35 0.22 0.0
Format Passed 31.04 8.68 31.75 39.19 14.76 38.78 52.36 14.61 35.59 52.42 19.37 35.63
Assertion 31.87 58.68 46.52 35.87 62.54 40.24 36.38 66.18 44.13 35.14 65.71 42.8
Compilation 33.15 31.18 21.02 20.55 20.51 20.61 7.72 16.95 19.72 9.19 12.66 21.36
Timeout 0.96 1.14 0.71 0.96 1.24 0.37 0.43 1.06 0.57 0.28 0.96 0.2
Runtime 2.99 0.33 0.0 3.43 0.96 0.0 3.11 1.2 0.0 2.97 1.3 0.0
RQ3: How the enriched test cases affect the performance of the LLMs under testing?
Motivation: The findings from a recent study (Liu et al., 2023) highlight the limitation of humaneval test cases which can not detect all of the incorrect samples. To enhance the precision and depth of our investigation, we use the translated test cases from evalplus for the evaluation of generated code in our study. By doing this, we aim to ensure a more thorough exploration of the errors generated by the models, enabling us to better uncover the scope of inaccuracies in their code generation among the three languages we consider.
Approach: To answer this research question, we follow the same procedure as we follow for HumanEval-X test cases. We use the same four models (Incoder1B, Incoder6B, CodeGen-Multi-2B, and CodeGen-Multi-6B) and three languages (Java, C++, and JavaScript), and the same metrics ($RP_{s}@k$, $RD_{s}@k$, and $RR_{s}@k$) to evaluate the performance.
Results: Using more test cases results in reduced model robustness compared to using only the original test cases. Table 11 and Table 12 present the results for worst-case and partial-case scenarios using EvalPlus test cases. Analyzing these results alongside Table 8 and Table 9 reveals a consistent performance decline across all cells. This suggests that EvalPlus test cases cause more failures in samples that previously passed the HumanEval-X test cases. For instance, Figure 5 shows a code example generated from a Back-Translation perturbed prompt that successfully passes a HumanEval-X test case but times out on a new EvalPlus test case.
Table 11: Performance summary of perturbations against code generation models using EvalPlus test cases and worst case passes (5 out of 5)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
@@
RP5@1$\uparrow$ 0.0 0.02 0.03 0.03 0.05 0.04 0.07 0.05 0.04 0.04 0.07 0.05
RD5@1$\downarrow$ 0.89 0.71 0.46 0.71 0.47 0.59 0.42 0.56 0.61 0.68 0.4 0.53
RR5@1$\downarrow$ 0.02 0.07 0.04 0.08 0.05 0.06 0.05 0.07 0.06 0.1 0.06 0.07
Syntax Nominal (Partial)$\uparrow$ 0.32 0.08 0.35 0.43 0.13 0.41 0.55 0.12 0.36 0.54 0.2 0.41
RP5@1$\uparrow$ 0.22 0.02 0.22 0.26 0.04 0.25 0.32 0.05 0.19 0.28 0.07 0.22
RD5@1$\downarrow$ 0.3 0.81 0.39 0.39 0.7 0.4 0.42 0.62 0.46 0.48 0.63 0.47
RR5@1$\downarrow$ 0.21 0.07 0.14 0.24 0.11 0.16 0.25 0.09 0.17 0.26 0.14 0.23
Format Nominal (Partial)$\uparrow$ 0.32 0.08 0.35 0.43 0.13 0.41 0.55 0.12 0.36 0.54 0.2 0.41
RP5@1$\uparrow$ 0.2 0.03 0.24 0.28 0.06 0.29 0.44 0.08 0.26 0.41 0.1 0.26
RD5@1$\downarrow$ 0.38 0.65 0.33 0.35 0.58 0.3 0.2 0.38 0.27 0.23 0.51 0.38
RR5@1$\downarrow$ 0.16 0.06 0.13 0.18 0.08 0.14 0.13 0.05 0.14 0.13 0.1 0.2
Table 12: Performance summary of perturbations using partial cases using EvalPlus test cases and partial passes (3 out of 5)
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-Multi CodeGen-6B-Multi
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS
@@
RP5@1$\uparrow$ 0.02 0.06 0.06 0.06 0.08 0.08 0.11 0.1 0.08 0.1 0.11 0.1
RD5@1$\downarrow$ -0.11 0.27 -0.02 0.35 0.12 0.16 0.08 0.16 0.09 0.26 0.09 0.12
RR5@1$\downarrow$ 0.02 0.06 0.04 0.05 0.04 0.04 0.04 0.06 0.05 0.06 0.04 0.05
Syntax Nominal (Partial)$\uparrow$ 0.32 0.08 0.35 0.43 0.13 0.41 0.55 0.12 0.36 0.54 0.2 0.41
RP5@1$\uparrow$ 0.26 0.05 0.26 0.35 0.09 0.3 0.38 0.08 0.27 0.38 0.12 0.31
RD5@1$\downarrow$ 0.17 0.42 0.25 0.2 0.36 0.28 0.31 0.34 0.26 0.29 0.39 0.25
RR5@1$\downarrow$ 0.21 0.08 0.1 0.2 0.13 0.13 0.21 0.09 0.14 0.18 0.12 0.17
Format Nominal (Partial)$\uparrow$ 0.32 0.08 0.35 0.43 0.13 0.41 0.55 0.12 0.36 0.54 0.2 0.41
RP5@1$\uparrow$ 0.31 0.05 0.29 0.38 0.11 0.4 0.51 0.11 0.34 0.53 0.16 0.33
RD5@1$\downarrow$ 0.01 0.33 0.18 0.12 0.2 0.04 0.08 0.09 0.06 0.02 0.16 0.2
RR5@1$\downarrow$ 0.11 0.04 0.09 0.14 0.07 0.06 0.09 0.05 0.1 0.05 0.05 0.14

### v4

@@ 5.1 RQ1: How robust are LLMs across different programming languages?
    A Larger model size does not always guarantee better robustness or pass rate across all programming languages. As shown in Table 8 and Table 10, in general, we can observe a robustness or pass rate improvement with the increase of model sizes. For instance, for CodeGen in the worst-case scenario, the $RD_{5}@1$ of JavaScript increases from 0.07 to 0.08 for 2B and 6B, respectively. On the contrary, for Java, we also observe that there is a drop of 0.02 of the $RD_{5}@1$. The results show that increasing the model size does not always guarantee better robustness and highlight the need for more comprehensive experiments on different programming languages.
  2. (2)
    CodeGen generally performs better than Incoder, while Magicoder and QwenCode further outperform both. As shown in Tables 8 and 10, CodeGen achieves higher nominal pass@k than Incoder at comparable scales (e.g., 0.16 vs. 0.11 on Java). This is consistent with their training data differences: CodeGen is trained on a much larger token corpus, whereas Incoder, although more multilingual, is heavily skewed toward Python and JavaScript, limiting its cross-language generalization. On the other hand, Magicoder-7B and QwenCode outperform both CodeGen and Incoder, but with different behaviors under perturbations. Magicoder shows higher robustness, maintaining stronger pass@k when prompts are partially truncated or reformatted. This can be attributed to its OSS-INSTRUCT training, which leverages diverse open-source code snippets to reduce synthetic data bias and improve instruction grounding. QwenCode, trained on massive curated multilingual corpora, excels in nominal accuracy but exhibits slightly larger drops under perturbation, suggesting its strong generation ability is more sensitive to prompt variation.
  3. (3)
    Stronger test cases expose more failures across all languages. From Tables 8 and 9, it is evident that the stronger test cases from EvalPlus reveal substantially more failures in both nominal and perturbed settings, which aligns with prior findings Liu et al. (2023). The same trend holds for the partial cases, as shown in Tables 10 and 11. However, the robustness drop remains roughly consistent regardless of the absolute performance levels.
5.2 RQ2: What are the major factors that contribute to the difference in the robustness of LLMs on different programming languages?
(a) Function Name Perturbation
(b) DocString Perturbation
(c) Syntax Perturbation
(d) Format Perturbation
Figure 7: The heatmaps show the impact score for each feature, calculated relative to the overall average.
Approach: To diagnose the underlying causes for the robustness disparities observed in RQ1, we employ a quantitative, feature-level quantitative analysis. Our approach is designed to identify which specific characteristics of the input prompts or the generated code are most strongly associated with failure in each language. This is achieved by calculating a weighted impact score for each feature and visualizing these scores in a comparative heatmap. The approach is executed in four sequential steps:
Step 1. Data Selection and Preparation. We aim to identify features correlated with the Robust Drop, which represents the performance degradation when models generate code from perturbed versus nominal prompts. Therefore, the dataset for this analysis is scoped to include only samples that were passed on the nominal (unperturbed) prompt. This scoping decision isolates failures induced by a perturbation, distinguishing them from problems the model failed on initially. Then our goal is to identify the features most strongly associated with the samples that subsequently fail on the perturbed prompt.
Feature Calculation. For each sample, we compute a comprehensive set of 18 features. These features are grouped into two categories:
Category 1. Input Perturbation Features: Quantify the magnitude of change in prompts.
  1. Feature 1-4. Token Dissimilarity (func_name_change, docstring_change, code_change, prompt_change): These features compute the percentage of token-level change between two code snippets (a nominal component $P_{n}$ and a perturbed component $P_{p}$). We first tokenize both snippets. We then calculate a similarity ratio by finding the total number of matching tokens within all matching subsequences, and relate this to the total number of tokens in both snippets, as defined in Equation 5.
    $\text{Dissimilarity \%}(P_{n},P_{p})=\left(1-\frac{2\cdot|\text{Matches}(P_{n},P_{p})|}{|P_{n}|+|P_{p}|}\right)\times 100$ (5)
    Here, $P_{n}$ and $P_{p}$ are the token sequences, and $|\text{Matches}|$ is the count of matching tokens from the sequence matcher.
  2. Feature 5. Comment Percentage (comment_percentage_p_prompt): Measures the ratio of tokens inside comments to the total number of tokens in the perturbed prompt.
    $\text{Comment \%}=\frac{|\text{Tokens in Comments}_{p}|}{|\text{Total Tokens in Prompt}_{p}|}$ (6)
  3. Feature 6. Scope Delimiter Count (scope_delimiter_count): Measures the quantity of syntax elements used to open code blocks and define scope (e.g., counting the opening brace { in C++/Java). This metric serves as a simple proxy for the code’s structural complexity, which may correlate to the model’s fragility.
2. Output and Problem Features: Quantify the characteristics of the canonical solutions and models generated output, which represents the complexity of the problem.
  1. Feature 7-9. Output Dissimilarity (generated_code_change, nominal_ code_change, perturbed_code_change): Measures the token dissimilarity (using Equation 5) between different versions of code to quantify changes in output structure and solution correctness.
    • generated_code_change: Dissimilarity between the model-generated code from nominal_prompt and perturbed_prompt. A high value indicates that the perturbation caused the model to generate a structurally different solution.
    • nominal_code_change: Dissimilarity between the human-written reference_code and the LLM’s generated code from nominal_prompt. This quantifies the changes in the nominal code generated by the LLM.
    • perturbed_code_change: Dissimilarity between the human-written reference_code and the generated code by the LLM from the perturbed prompt. This quantifies the changes in the perturbed code generated by the LLM.
  2. Feature 10-12. Lines of Code (LOC) (nominal_LOC, perturbed_LOC, reference_LOC): A count of the total lines of code for the nominal generated code, perturbed generated code, and reference solution, respectively.
  3. Feature 13-15. Token Count (nominal_tokens, perturbed_tokens, reference_tokens): A count of the total tokens for the nominal generated code, perturbed generated code, and reference solution, respectively.
  4. Feature 16-18. Cyclomatic Complexity (nominal_complexity, pertur- bed_complexity, reference_complexity): Measures the structural complexity of the code (e.g., McCabe’s complexity) by counting linearly independent paths. This is calculated for the nominal generated code, perturbed generated code, and reference solution.
Statistical Calculation of the Impact Score To create a fair, language-to-language comparison, we calculate an Impact Score for every feature. This score quantifies the relationship between each feature and the robustness failures, addressing a key challenge: a feature that is highly associated with failure in a language that less-frequently fails is less impactful overall than a feature associated with failure in a language that frequently fails. The Impact Score ($I$) for a given feature ($F$) and language ($L$) is the product of the feature’s Z-Score ($Z$) and the language’s Failure Rate (FR).
The calculation follows these steps for each language $L$:
  1. (1)
    Calculate Failure Rate ($\text{FR}_{L}$): The proportion of samples that failed on the perturbed prompt.
    $\text{FR}_{L}=\frac{|\text{Failed Samples}_{L}|}{|\text{Total Samples}_{L}|}$ (7)
  2. (2)
    Establish Language-Specific Statistical Norms: For each feature $F$, we calculate the mean ($\mu_{F,L}$) and standard deviation ($\sigma_{F,L}$) using all samples (passed and failed) for that language $L$.
  3. (3)
    Profile Failed Runs: We calculate the mean of the failed samples ($\mu_{F,L}^{\text{fail}}$) for that feature $F$ and language $L$.
  4. (4)
    Calculate Z-Score ($Z_{F,L}$): We standardize the failed-run average against its language-specific statistical norms. This use of Z-score normalization is a standard method for data preprocessing. It is necessary here to transform features with heterogeneous units and scales (e.g., nominal_LOC versus func_name_change) onto a common, dimensionless scale. This allows us to compare the relative unusualness of each feature’s value, as the resulting Z-score represents the deviation from the language-specific mean in units of standard deviations. A high positive Z-Score means failed runs are strongly associated with high values of that feature.
    $Z_{F,L}=\frac{\mu_{F,L}^{\text{fail}}-\mu_{F,L}}{\sigma_{F,L}}$ (8)
  5. (5)
    Calculate Impact Score ($I_{F,L}$): We weigh the Z-Score by the language’s overall Failure Rate.
    $I_{F,L}=Z_{F,L}\times\text{FR}_{L}$ (9)
Visualization. The final output is a matrix of Impact Scores, with languages as rows and features as columns. This matrix is visualized as a 1D-stacked heatmap for each language. A diverging colormap (e.g., blue-white-red) is centered at zero and clipped to a fixed scale (e.g., -0.2 to +0.2) to ensure all heatmaps are directly comparable. This final visualization directly highlights which features are the most significant drivers of robustness failures for each language. 7
Result. Figure 7 illustrates the correlation between features and robustness drop across programming languages. In each plot of the figure, the y-axis represents the programming languages, while the x-axis represents the evaluated features. Each cell displays an impact score, calculated using Equations 8 and 9. A higher positive value indicates a stronger positive correlation between the feature and the language, whereas a negative value suggests a negative correlation. A value of zero indicates no relationship. Each subplot corresponds to one perturbation type (Function Name, DocString, Syntax, and Format). Rows represent programming languages (C++, Java, and JavaScript), while columns correspond to the features. The color and magnitude of each cell indicate the Impact Score, which reflects the strength and direction of correlation between a given feature and the robustness drop for that language.
Overall, the heatmaps reveal language-specific patterns of sensitivity under different perturbations. JavaScript consistently shows the highest positive correlations across multiple features, aligning with its higher robustness drop reported in RQ1. In contrast, Java and C++ demonstrate more stable patterns, with lower overall correlation magnitudes and fewer dominant features.
Across the four perturbation types, the overall trend shows that Java tends to exhibit the strongest feature impact relationships under function name and docString perturbations, where several features (e.g., generated_code_change, perturbed_code_change) consistently reach impact scores around 0.10-0.15. In contrast, this pattern does not hold under syntax perturbations, where Java’s impact values drop noticeably, and most features fall close to zero, making Java the least sensitive language in this specific condition. C++ shows moderate but steady correlations across most perturbations, often around 0.05-0.11, without strong spikes, indicating a stable but weaker relationship between features and robustness drop. JavaScript generally exhibits low or near-zero correlations across all perturbation types, occasionally showing small positive effects (e.g., perturbed_tokens, reference_tokens) but overall demonstrating the flattest patterns. Taken together, the results indicate that Java is feature-sensitive under perturbations in function name and docString, while syntax and formatting perturbations suppress its feature–impact relationships, making correlations there more diffuse, whereas C++ remains moderate and steady, and JavaScript remains consistently low across all cases.
Summary of Findings. Across perturbations, several consistent trends emerge:
  1. (1)
    Changes in the DocString have a significant impact on the robustness drop for Java and JavaScript, and a moderate impact for C++. From the DocString perturbation, we observe that even small modifications to the DocString, caused indirectly by other name changes, can also lead to notable performance variations.
  2. (2)
    For the generated code from perturbed prompts, we observe substantial deviations compared to the code generated from nominal prompts and the original human-written solutions.
  3. (3)
    For function name and DocString perturbations, many features strongly affect robustness in Java and JavaScript but have a smaller influence on C++. In contrast, C++ is more affected by syntax-level perturbations than the other languages.
@@
Table 8: Performance summary of perturbations against code generation models using EvalPlus test cases and worst case passes (5 out of 5). Up arrows indicate the extent of performance improvement, while down arrows indicate the extent of performance drop.
HumanEval-X Model Incoder-1B Incoder-6B CodeGen-2B-multi CodeGen-6B-multi Magicoder7B QwenCode
Perturbation Metric Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS Java CPP JS
Nominal RP5@1 0.02 0.08 0.07 0.11 0.1 0.12 0.14 0.12 0.11 0.15 0.13 0.14 0.63 0.55 0.51 0.55 0.57 0.6
Partial RP5@1 0.32 0.09 0.37 0.45 0.14 0.44 0.56 0.13 0.39 0.55 0.2 0.44 0.86 0.61 0.74 0.77 0.63 0.68
@@
[…]